# Scalable Multi-Stage Influence Function for Large Language Models via Eigenvalue-Corrected Kronecker-Factored Parameterization

### v1

@@ Abstract
In this paper, we propose the multi-stage influence function to attribute the downstream predictions of fine-tuned LLMs to pre-training data under the full-parameter fine-tuning paradigm. To enhance the efficiency and practicality of our multi-stage influence function, we leverage Eigenvalue-corrected Kronecker-Factored (EK-FAC) parameterization for efficient approximation. Empirical results validate the superior scalability of EK-FAC approximation and the effectiveness of our multi-stage influence function. Additionally, case studies on a real-world LLM, dolly-v2-3b, demonstrate its interpretive power, with exemplars illustrating insights provided by multi-stage influence estimates. [^1] Our code is public at https://github.com/colored-dye/multi_stage_influence_function
@@ 4.1 Multi-Stage Influence Function
$\boldsymbol{\theta}^{\text{ft}}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}_{\text{ft}}(\boldsymbol{\theta})+\frac{\alpha}{2}||\boldsymbol{\theta}-\boldsymbol{\theta}^{\text{pt}}||_{2}^{2},$ (11)
@@ 5 Experiments
This section presents a series of experiments to evaluate our proposed method. First, we assess the scalability of single-stage IFs approximated using EK-FAC compared to various TDA methods in terms of both estimation accuracy and wall-clock runtime on language modeling tasks (Section 5.1). Next, we investigate the validity of the underlying assumption of our multi-stage IF (Section 5.2). Subsequently, we evaluate the effectiveness of the multi-stage IF on a factual knowledge retrieval task (Section 5.3). Finally, we showcase a qualitative case study using the multi-stage IF on an instruction-following LLM (Section 5.4).
@@ 5.1 Scalability Validation of EK-FAC
Spearman $\rho$uparrow Overhead Time $\downarrow$ Pair-wise Time $\downarrow$
@@ Results.
    Arnoldi yields poor estimates at $\sim$120$\times$ the pair-wise compute cost of EK-FAC.
@@
Figure 2: Distribution of $||\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}$ and $||\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}/||\boldsymbol{\theta}^{\text{pt}}||_{2}.$
@@ 5.2 Euclidean Proximity in Practice
The results, presented in Figure 2, indicate that the distance between fine-tuned and pre-trained weights ($||\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}$) accounts for no more than $8\%$ of the L2 norm of the pre-trained weights, $||\boldsymbol{\theta}^{\text{pt}}||_{2}$. Specifically, for the BLOOM-560m and BLOOMZ-560m pair, $||\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}^{2}=368.7$, while for the Pythia-2.8b and dolly-v2-3b pair, $||\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}^{2}=94.5$.

### v2

@@ Abstract
In this paper, we propose the multi-stage influence function to attribute the downstream predictions of fine-tuned LLMs to pre-training data under the full-parameter fine-tuning paradigm. To enhance the efficiency and practicality of our multi-stage influence function, we leverage Eigenvalue-corrected Kronecker-Factored (EK-FAC) parameterization for efficient approximation. Empirical results validate the superior scalability of EK-FAC approximation and the effectiveness of our multi-stage influence function. Additionally, case studies on a real-world LLM, dolly-v2-3b, demonstrate its interpretive power, with exemplars illustrating insights provided by multi-stage influence estimates.
@@ 4.1 Multi-Stage Influence Function
[…]
@@ 5 Experiments
This section presents a series of experiments to evaluate our proposed method. First, we assess the scalability of single-stage IFs approximated using EK-FAC compared to various TDA methods in terms of both estimation accuracy and wall-clock runtime on language modeling tasks (Section 5.1). Next, we investigate the validity of the underlying assumption of our multi-stage IF (Section 5.2). Subsequently, we evaluate the effectiveness of the multi-stage IF on a factual knowledge retrieval task (Section 5.3). Finally, we showcase a qualitative case study using the multi-stage IF on an instruction-following LLM (Section 5.4). Our code is public at https://github.com/colored-dye/multi_stage_influence_function.
@@ 5.1 Scalability Validation of EK-FAC
Spearman $\rho$ $\uparrow$ Overhead Time $\downarrow$ Pair-wise Time $\downarrow$
@@ Results.
[…]
@@
[…]
@@ 5.2 Euclidean Proximity in Practice
The results, presented in Figure 2, indicate that the distance between fine-tuned and pre-trained weights ($\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}$) accounts for no more than $8\%$ of the L2 norm of the pre-trained weights, $\boldsymbol{\theta}^{\text{pt}}||_{2}$. Specifically, for the BLOOM-560m and BLOOMZ-560m pair, $\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}^{2}=368.7$, while for the Pythia-2.8b and dolly-v2-3b pair, $\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pt}}||_{2}^{2}=94.5$.